In rhombus ABCD, the measure, in inches, of AB is 3x + 2 and BC is x + 12. Find the number of inches in the length of DC.
Answer by nyc_function(2733)

A rhombus has equal measure all around.
This means we can equate the given expressions and solve for x as step one.
3x + 2 = x + 12
3x - x = -2 + 12
2x = 10
x = 10/2
x = 5
Now, side AB = side DC.
AB was given to be 3x + 2.
Let x = 5
AB = 3(5) + 2
AB = 15 + 2
AB = 17 inches
If AB = DC, then DC = 17 inches